# Fractional diffusion as the limit of a short range potential Rayleigh gas

### v1

@@ 4.3. The set of re-collisions
$\leq C(M(\varepsilon))^{2}\left(\frac{\tilde{\eta}V(\varepsilon)\varepsilon}{\tilde{\delta}^{2}}+\tilde{\delta}+\frac{TV(\varepsilon)}{\tilde{\eta}}\left(\frac{\varepsilon}{\tilde{\eta}}\right)^{2}\right).$ (4.4)
@@
The main strategy here is to identify variables enforcing the re-collision and then estimate the volume of possible velocities or collision parameters. We will consider various cases. Some estimates are easier than the qualitative estimates in [MST18] because we are on the domain $\mathds{R}^{3}$ such that there will be no re-collisions due to periodic boundary conditions on the torus. As we are estimating events for the idealised distribution, we consider idealised trajectories with instantaneous collision events for point particles, but that come close within a distance $\varepsilon$.
@@
where $x(s)=x_{0}+t_{1}v_{0}+(t_{2}-t_{1})v(t_{1})+...+(s-t_{j})v(t_{j})$ is the position of the tagged particle and $x_{k}(s)=x_{k}(t_{k})+\varepsilon\nu+(s-(t_{k}))\bar{v}_{k}$ is the position of the re-colliding background particle. Then (4.5) is equivalent to
$Y+(s-t_{j})v(t_{j})+\varepsilon\tilde{\nu}=(s-t_{k})\bar{v}_{k},$
where $Y:=(t_{k}-t_{k-1})v(t_{k-1})+...+(t_{j}-t_{j-1})v(t_{j-1})$. This can be solved for the terminal velocity of the tagged particle
$v(t_{j})=\frac{-\varepsilon\tilde{\nu}}{s-t_{j}}+\frac{\tilde{Y}-s\bar{v}_{k}}{t_{j}-s},$
where $\tilde{Y}:=Y+t_{k}\bar{v}_{k}.$ By varying $s\in(t_{j},T]$ and fixing $\bar{v}_{k}$ as above, then $v(t_{j})$ is contained in the "cylinder" with varying radius $\frac{\varepsilon}{s-t_{j}}$ around the curve defined by
$r(s):=\frac{\tilde{Y}(t)-s\bar{v}_{k}}{t_{j}-s},$ (4.6)
for $1\leq k<j$. We split the interval $(t_{j},T]$ in two parts. The first part is $(t_{j},t_{j}+\tilde{\eta})$ which covers re-collision shortly after the last collision with $\tilde{\eta}>0$ sufficiently small to be chosen later and the bulk part $[t_{j}+\tilde{\eta},T]$.
Case 1: $s\in[t_{j}+\tilde{\eta},T]$
We restrict the interval to $s\in[t_{j}+\tilde{\eta},T]$ and consider the differentiable curve $r$ in (4.6) for $s>t_{j}$. We estimate its length by calculating
$\frac{\mathrm{d}}{\mathrm{d}s}r(s)=\frac{\tilde{Y}(t)-s\bar{v}_{k}}{(t_{j}-s)^{2}}-\frac{\bar{v}_{k}}{t_{j}-s}.$ (4.7)
We observe that
$|\tilde{Y}|$ $\leq(t_{k}-t_{k-1})|v(t_{k-1})|+...+(t_{j}-t_{j-1})|v(t_{j-1})|+t_{k}|\bar{v}_{k}|$
$\leq(t_{k}-t_{k-1})V(\varepsilon)+...+(t_{j}-t_{j-1})V(\varepsilon)+t_{k}V(\varepsilon)$
$=(t_{j}-t_{k-1})V(\varepsilon)+t_{k}V(\varepsilon)\leq TV(\varepsilon)+TV(\varepsilon).$
Therefore,
$|\tilde{Y}|\leq 2TV(\varepsilon),\ \varepsilon>0.$ (4.8)
Thus using (4.8) in (4.7),
$\left|\frac{\mathrm{d}}{\mathrm{d}s}r(s)\right|$ $\leq\frac{|\tilde{Y}(t)-s\bar{v}_{k}|}{(t_{j}-s)^{2}}+\frac{|\bar{v}_{k}|}{|t_{j}-s|}\leq\frac{2TV(\varepsilon)+TV(\varepsilon)}{(t_{j}-s)^{2}}+\frac{V(\varepsilon)}{|t_{j}-s|}.$
Therefore the length of the curve for $s\in[t_{j}+\tilde{\eta},T]$ is bounded by
$\int_{t_{j}+\tilde{\eta}}^{T}\left|\frac{\mathrm{d}}{\mathrm{d}s}r(s)\right|\,\mathrm{d}s$ $\leq\int_{t_{j}+\tilde{\eta}}^{T}\frac{3TV(\varepsilon)}{(t_{j}-s)^{2}}\,\mathrm{d}s+\int_{t_{j}+\tilde{\eta}}^{T}\frac{V(\varepsilon)}{|t_{j}-s|}\,\mathrm{d}s.$
Then evaluating the first integral
$\int_{t_{j}+\tilde{\eta}}^{T}\frac{3TV(\varepsilon)}{(t_{j}-s)^{2}}\,\mathrm{d}s$ $=3TV(\varepsilon)\left[-\frac{1}{T-t_{j}}+\frac{1}{\tilde{\eta}}\right]\leq\frac{3TV(\varepsilon)}{\tilde{\eta}}.$
The second integral is
$\int_{t_{j}+\tilde{\eta}}^{T}\frac{V(\varepsilon)}{s-t_{j}}\,\mathrm{d}s$ $=V(\varepsilon)[\log(T-t_{j})-\log(\tilde{\eta})]$
$\leq V(\varepsilon)[\log T-\log\tilde{\eta}]=V(\varepsilon)\log\left(\frac{T}{\tilde{\eta}}\right).$
Here, we used the monotonicity of $\log$. Hence overall, using $\log(T)<T$
$\mathrm{length}(r_{|s\in[t_{j}+\tilde{\eta},T]})\leq 3TV(\varepsilon)\left(\frac{1}{\tilde{\eta}}\right)+V(\varepsilon)\log\left(\frac{T}{\tilde{\eta}}\right)$ $<4TV(\varepsilon)\left(\frac{1}{\tilde{\eta}}\right)=:\hat{C}(\tilde{\eta}).$
Therefore, there exists $C>0$ such that the volume of re-colliding velocities for $1\leq k<j$ and $2\leq j\leq n(\Phi)$ and $s\in[t_{j}+\tilde{\eta},T]$ has the bound
$\text{vol}(v(t_{j}))\leq C\hat{C}(\tilde{\eta})\left(\frac{\varepsilon}{\tilde{\eta}}\right)^{2}.$ (4.9)
Case 2: $s\in(t_{j},t_{j}+\tilde{\eta})$
@@
The factor $s-t_{k}$ can be easily bounded from below by some $\tilde{\delta}>0$ with a probability $1-\tilde{\delta}$ as $k<j$. Then the overall volume of suitable $v_{k}$ and $\nu_{k}$ can be estimated
$\mathrm{vol}(v_{k},\nu_{k})\leq C\frac{1}{\tilde{\delta}^{2}}\tilde{\eta}V(\varepsilon)\varepsilon.$ (4.14)
@@
[…]
@@
$\leq C\sum_{j=2}^{n(\Phi)}M(\varepsilon)\left(\frac{\tilde{\eta}V(\varepsilon)\varepsilon}{\tilde{\delta}^{2}}+\tilde{\delta}+\frac{TV(\varepsilon)}{\tilde{\eta}}\left(\frac{\varepsilon}{\tilde{\eta}}\right)^{2}\right)$
$\leq C(M(\varepsilon))^{2}\left(\frac{\tilde{\eta}V(\varepsilon)\varepsilon}{\tilde{\delta}^{2}}+\tilde{\delta}+\frac{TV(\varepsilon)}{\tilde{\eta}}\left(\frac{\varepsilon}{\tilde{\eta}}\right)^{2}\right),$
@@ 4.4. Initial overlap
Let $\Phi\in\mathcal{MT}$ such that $\Phi=((x_{0},v_{0}),(t_{1},\nu_{1},v_{1}),...,(t_{n},\nu_{n},v_{n}))$, where $(x_{0},v_{0})\in\mathds{R}^{3}\times\mathds{R}^{3}$, $t_{i}\in[0,T]$, $\nu_{i}\in\mathbb{S}^{2}$, $v_{i}\in\mathds{R}^{3}$, for $i=1,...,n$. The condition for initial overlap of supports of the potentials is
@@
$|Q^{-}_{t}(\Phi)|=\left|\int_{\mathds{R}^{3}}\int_{\mathbb{S}^{2}}g_{0}(\bar{v})B(v_{0}-\bar{v},\omega)\,\mathrm{d}{\omega}\,\mathrm{d}\bar{v}\right|\leq C,$
due to the form of $B$. Hence the probability to have two collisions is bounded by
@@ 4.5. Quantitative estimates for bad trees
For any decreasing functions $V,M:(0,\infty)\to[0,\infty)$ such that $\lim_{\varepsilon\to 0}V(\varepsilon)=\lim_{\varepsilon\to 0}M(\varepsilon)=\infty$ and for any $q>4$, $t\in[0,T]$, $\delta>0$, $\tilde{\delta}>0$, $\tilde{\eta}>0$ and $n\in(2,5]$, the probability of the set of bad trees has the following error estimate
$P_{t}(\mathcal{MT}\setminus\mathcal{G}(\varepsilon))$ $\leq\textnormal{Ov}(\varepsilon)+\textnormal{Rec}(\varepsilon)+\textnormal{Hi}(\varepsilon)+\textnormal{Vel}(\varepsilon),$ (4.19)
@@
    $\textnormal{Rec}(\varepsilon)\leq C(M(\varepsilon))^{2}\left(\frac{\tilde{\eta}V(\varepsilon)\varepsilon}{\tilde{\delta}^{2}}+\tilde{\delta}+\frac{TV(\varepsilon)}{\tilde{\eta}}\big{(}\frac{\varepsilon}{\tilde{\eta}}\big{)}^{2}\right)$ is due to re-collisions,
@@
    $\textnormal{Hi}(\varepsilon)\leq\frac{1}{M(\varepsilon)}(1+Cct)$ is due to the unbounded number of collisions and
@@
$+P_{t}\big{(}\mathcal{V}(\Phi)>V(\varepsilon)\ \mathrm{with}\ n(\Phi)\leq M(\varepsilon)\big{)}.$
We observe that the third term can be written as
@@
for all $t\geq 0$ and $n\in(2,5]$. Also, by Proposition 4.4 we have
$\leq C(M(\varepsilon))^{2}\left(\frac{\tilde{\eta}V(\varepsilon)\varepsilon}{\tilde{\delta}^{2}}+\tilde{\delta}+\frac{TV(\varepsilon)}{\tilde{\eta}}\left(\frac{\varepsilon}{\tilde{\eta}}\right)^{2}\right).$ (4.22)
@@
by using Proposition 4.2 in the last inequality, since $\tilde{M}_{f}(t)\leq D_{f}(t)$. Therefore, by summing (4.15), (4.21), (4.4) and (4.23) we obtain
@@
$\quad+C(M(\varepsilon))^{2}\left(\frac{\tilde{\eta}V(\varepsilon)\varepsilon}{\tilde{\delta}^{2}}+\tilde{\delta}+\frac{TV(\varepsilon)}{\tilde{\eta}}\left(\frac{\varepsilon}{\tilde{\eta}}\right)^{2}\right)$
$\quad+\frac{M_{f_{0}}}{V(\varepsilon)}+(C_{g}+\tilde{C}_{q,n}ct)\frac{M(\varepsilon)}{V(\varepsilon)},\ t\in[0,T]\ \textrm{and}\ n\in(2,5].$
@@ 4.6. The Empirical Distribution
$\mathds{1}^{\varepsilon}_{t}[\Phi](\bar{x},\bar{v}):=\begin{cases}1&\text{if\ for\ all\ }s\in(0,t),\ |x(s)-(\bar{x}+s\bar{v})|>\varepsilon\text{ or }\\ &\quad\text{if\ for\ all\ }s\in(0,t)\text{ with }|x(s)-(\bar{x}+s\bar{v})|\leq\varepsilon:\beta(\bar{x}+s\bar{v},\bar{v},s)=0\\ 0&\text{else}.\end{cases}$ (4.24)
@@
$\hat{\mathcal{Q}}^{+}_{t}[\hat{P_{t}}](\Phi):=\begin{cases}\delta(t-\tau)\hat{P_{t}}(\bar{\Phi})g_{0}(v^{\prime})B(v(\tau^{-})-v^{\prime},\omega)&\text{if}\ n(\Phi)\geq 1\\ 0&\text{if}\ n(\Phi)=0.\end{cases}$
@@
$\hat{\mathcal{Q}}^{-}_{t}[\hat{P_{t}}](\Phi):=\mathds{1}_{t\geq\tau(\Phi)+T^{*}(\Phi)}\hat{P_{t}}(\Phi)\int_{\mathbb{S}^{2}}\int_{\mathds{R}^{3}}g_{0}(\bar{v})B(v(\tau)-\bar{v},\omega)\mathds{1}^{\varepsilon}_{t}[\Phi](x(t)+\varepsilon\omega,\bar{v})\,\mathrm{d}\bar{v}\,\mathrm{d}\omega.$
@@
$\begin{cases}\partial_{t}{\hat{P_{t}}}(\Phi)&=c\hat{\mathcal{Q}_{t}}[\hat{P_{t}}](\Phi)\\ \hat{P_{0}}(\Phi)&=f_{0}(x_{0},v_{0})\mathds{1}_{n(\Phi)=0}.\end{cases}$ (4.25)
This is direct adaption of [MST18, Theorem 4.6.] and [Sto17] for the version of Rayleigh in (1.3). The same approach was also used in [MS18, ET17, Egg18]. There are two substantial simplification compared to [MST18]: firstly due to the Poisson set-up we do not have terms due to conditioning on the history or due to a reduction of available particles and secondly due to our definition of the collision status in (1.2) any initial overlap can be ignored.
@@
$\hat{\mathcal{Q}}^{-}_{t}[\hat{P_{t}}](\Phi)$ $=\hat{L}_{t}(\Phi)\hat{P_{t}}(\Phi),$
$\mathcal{Q}^{-}_{t}[P_{t}](\Phi)$ $=L_{t}(\Phi)P_{t}(\Phi)$
@@
Then for a fixed $\Phi$ and $t>\tau(\Phi)$ we obtain a scalar ordinary differential inequality for $\hat{P_{t}}(\Phi)-P_{t}(\Phi)$
$\partial_{t}(\hat{P_{t}}(\Phi)-P_{t}(\Phi))$ $=-c\hat{L}_{t}(\Phi)\hat{P_{t}}(\Phi)+cL_{t}(\Phi)P_{t}(\Phi)$
$=-cL_{t}(\Phi)(\hat{P_{t}}(\Phi)-P_{t}(\Phi))+c[L_{t}(\Phi)-\hat{L}_{t}(\Phi)]\hat{P_{t}}(\Phi)$
$\geq-cL_{t}(\Phi)(\hat{P_{t}}(\Phi)-P_{t}(\Phi)).$
@@
$\hat{P_{t}}(\Phi)-P_{t}(\Phi)\geq\exp\left(-c\int_{\tau}^{t}L_{s}(\Phi)\mathrm{d}s\right)(\hat{P_{\tau}}(\Phi)-P_{\tau}(\Phi)).$ (4.28)
So we obtain (4.26) if $\hat{P_{\tau}}(\Phi)-P_{\tau}(\Phi)\geq 0$. The $n(\Phi)=0$ holds for $\tau=0$ due to the initial conditions in (2.1) and (4.25). For all other good trees this follows by induction over $n(\Phi)$ with (4.27) as the gain operator $\mathcal{Q}^{+}_{t}[P](\Phi)$ is monotone in $P$. ∎
@@
$\quad+C(M(\varepsilon))^{2}\left(\frac{\tilde{\eta}V(\varepsilon)\varepsilon}{\tilde{\delta}^{2}}+\tilde{\delta}+\frac{TV(\varepsilon)}{\tilde{\eta}}\left(\frac{\varepsilon}{\tilde{\eta}}\right)^{2}\right)$
$\quad+\frac{M_{f_{0}}}{V(\varepsilon)}+(C_{g}+Cct)\frac{M(\varepsilon)}{V(\varepsilon)}.$ (4.29)

### v2

@@ 4.3. The set of re-collisions
$P_{t}\big($ $\text{ re-collisions with}\ n(\Phi)\leq M(\varepsilon)\ and\ \mathcal{V}(\Phi)<V(\varepsilon)\big)$
@@
The main strategy here is to identify variables enforcing the re-collision and then estimate the volume of possible velocities of the background particles or collision parameters. It is enough to consider their volumes as we have uniform $L^{\infty}$ bounds on their densities and on the collision kernel $\mathcal{B}$. We will consider various cases. Some estimates are easier than the qualitative estimates in [MST18] because we are on the domain $\mathds{R}^{3}$ such that there will be no re-collisions due to periodic boundary conditions on the torus. As we are estimating events for the idealised distribution, we consider idealised trajectories with instantaneous collision events for point particles, but that come close within a distance $\varepsilon$.
@@
[…]
@@
The factor $s-t_{k}$ can be easily bounded from below by some $\tilde{\delta}>0$ with a probability $1-\tilde{\delta}$ for the case k=j1k=j-1 and with the same probability by $\sqrt{\tilde{\delta}}>0$ for k<j1k<j-1. Then the overall volume of suitable $v_{k}$ and $\nu_{k}$ can be estimated for k=j1k=j-1
$\mathrm{vol}(v_{k},\nu_{k})\leq C\frac{1}{\tilde{\delta}^{2}}\tilde{}\eta V(\varepsilon)\varepsilon$ (4.10)
and for the other $k$ by
$\mathrm{vol}(v_{k},\nu_{k})\leq C\frac{1}{\tilde{\delta}}\tilde{}\eta V(\varepsilon)\varepsilon$ (4.11)
@@
Case 2: $s\in[t_{j}+\tilde{\eta},T]$
We restrict the interval to $s\in[t_{j}+\tilde{\eta},T]$. Then (4.5) is equivalent to
$Y+sv(t_{j})+\varepsilon\tilde{\nu}=(s-t_{k})\bar{v}_{k},$
where $Y:=(t_{k}-t_{k-1})v(t_{k-1})+...+(t_{j}-t_{j-1})v(t_{j-1})-t_{j}v(t_{j})$. This can be solved for the terminal velocity of the recolliding particle and projecting onto $\nu_{k}^{\perp}$ again we obtain constraints in the "cylinder" with varying radius $\frac{\varepsilon}{s-t_{k}}$ around the curve defined by
$r(s):=\frac{Y+sv(t_{j})}{s-t_{k}},$ (4.12)
for $1\leq k<j$. We observe that $|Y+sv(t_{j})|\leq(s-t_{k})V(\varepsilon)$. Then we estimate its length by calculating
$\frac{\mathrm{d}}{\mathrm{d}s}r(s)=-\frac{Y+sv(t_{j})}{(s-t_{k})^{2}}+\frac{v(t_{j})}{s-t_{k}}.$ (4.13)
Therefore the $2d$ volume around the curve for $s\in[t_{j}+\tilde{\eta},T]$ is bounded by
$\int_{t_{j}+\tilde{\eta}}^{T}\left|\frac{\mathrm{d}}{\mathrm{d}s}r(s)\right|\frac{\varepsilon}{s-t_{k}}\,\mathrm{d}s$ $\leq\int_{t_{j}+\tilde{\eta}}^{T}\frac{(s-t_{k})V(\varepsilon)\varepsilon}{(s-t_{k})^{3}}\,\mathrm{d}s+\int_{t_{j}+\tilde{\eta}}^{T}\frac{V(\varepsilon)\varepsilon}{(s-t_{k})^{2}}\,\mathrm{d}s$
$\leq\int_{t_{j}+\tilde{\eta}}^{T}\frac{V(\varepsilon)\varepsilon}{(s-t_{k})^{2}}\,\mathrm{d}s+\int_{t_{j}+\tilde{\eta}}^{T}\frac{V(\varepsilon)\varepsilon}{(s-t_{k})^{2}}\,\mathrm{d}s\leq\frac{CV(\varepsilon)\varepsilon}{\tilde{\eta}+(t_{j}-t_{k})},$
where again $t_{j}-t_{j-1}\geq\tilde{\delta}$ and $t_{j}-t_{k}\geq\sqrt{\tilde{\delta}}$ for k<j1k<j-1.
@@
$P_{t}\big($ $\text{ re-collisions with}\ n(\Phi)\leq M(\varepsilon)\ and\ \mathcal{V}(\Phi)<V(\varepsilon)\big)$
@@
[…]
@@ 4.4. Initial overlap
Let $\Phi\in\mathcal{MT}$ such that $\Phi=((x_{0},v_{0}),(t_{1},\nu_{1},v_{1}),...,(t_{n},\nu_{n},v_{n}))$, where $(x_{0},v_{0})\in\mathds{R}^{3}\times\mathds{R}^{3}$, $t_{i}\in[0,T]$, $\nu_{i}\in\mathbb{S}^{2}$, $v_{i}\in\mathds{R}^{3}$, for i=1,,ni=1,...,n. The condition for initial overlap of supports of the potentials is
@@
$|Q^{-}_{t}(\Phi)|=\left|\int_{\mathds{R}^{3}}\int_{\mathbb{S}^{2}}g_{0}(\bar{v})\mathcal{B}(v_{0}-\bar{v},\omega)\,\mathrm{d}{\omega}\,\mathrm{d}\bar{v}\right|\leq C,$
due to the form of $\mathcal{B}$. Hence the probability to have two collisions is bounded by
@@
4.5. The set of binary collisions
Collisions of multiple background particles with the tagged particle are rare even for the empirical particle dynamics. According to [GSRT13][Prop. 8.2.1], individual collision times are bounded by $\tau^{*}\varepsilon$ with $\tau_{*}$ uniformly bounded on compact sets of $\mathcal{E}_{0}$ in $(0,\infty)$. On the level of the idealised distribution we can quantify estimates in terms of the minimal time interval $I(\varepsilon)$.
Lemma 4.6.
Let $I$ be a monotone and continuous function $I:[0,\infty)\to[0,\infty)$ with $\lim_{\varepsilon\to 0}I(\varepsilon)=0$. Then
$P_{t}\bigl(\{\Phi\in\mathcal{MT}:n(\Phi)\leq M(\varepsilon),\Phi\mbox{ is re-collision-free and }\vartheta(\Phi)\geq I(\varepsilon)\}\bigr)$ $\leq CM(\varepsilon)I(\varepsilon).$ (4.18)
Proof.
The rate of collisions is bounded, hence the probability to have a collision within the time $\vartheta(\Phi)$ is proportional to that time. As there are at most $M(\varepsilon)$ collisions in $\Phi$ and as each collision is independent we obtain the bound. ∎
4.6. Collisions with small relative velocities
To obtain uniform bounds on the collision times of the empiric realisation of a tree $\Phi$, we require also lower bounds on the relative velocities in each collision. The requirements rules out a small ball in velocity space around the pre-collision velocity $v(t_{j}^{-})$ of the tagged particle.
Lemma 4.7.
Let $J$ be a monotone and continuous function $J:[0,\infty)\to[0,\infty)$ with $\lim_{\varepsilon\to 0}J(\varepsilon)=0$. Then
$P_{t}\left(\{\Phi\in\mathcal{MT}:n(\Phi)\leq M(\varepsilon),\Phi\mbox{ is re-collision-free and }\iota(\Phi)<J(\varepsilon)\}\right)$ $\leq CM(\varepsilon)J^{3}(\varepsilon).$ (4.19)
Proof.
The probability to have a collision such that $|v_{j}-v(t_{j}^{-})|<J(\varepsilon)$, means it is enough to estimate
$\int_{|v-v(t_{j}^{-})|\leq J(\varepsilon)}g_{0}(v)\,\mathrm{d}v\leq C\left(J(\varepsilon)\right)^{3}$
for each collision. There are at most $M(\varepsilon)$ collisions in $\Phi$, which yields the bound. ∎
@@ 4.7. Quantitative estimates for bad trees
For any decreasing functions $V,M:(0,\infty)\to[0,\infty)$ such that $\lim_{\varepsilon\to 0}V(\varepsilon)=\lim_{\varepsilon\to 0}M(\varepsilon)=\infty$ and for any $q>4$, $t\in[0,T]$, $\delta>0$, $\tilde{\delta}>0$, $\tilde{\eta}>0$ and $n\in(3,5]$, the probability of the set of bad trees has the following error estimate
$P_{t}(\mathcal{MT}\setminus\mathcal{G}(\varepsilon))$ $\leq\textnormal{Ov}(\varepsilon)+\textnormal{Rec}(\varepsilon)+\textnormal{Hi}(\varepsilon)+\textnormal{Vel}(\varepsilon)+\textnormal{Tri}(\varepsilon)+\textnormal{Slow}(\varepsilon),$ (4.20)
@@
[…]
@@
    $\textnormal{Hi}(\varepsilon)\leq\frac{1}{M(\varepsilon)}(1+Cct)$ is due to the unbounded number of collisions,
@@
  5. (5)
    $\textnormal{Tri}(\varepsilon)\leq CM(\varepsilon)I(\varepsilon)$ is to rule small minimal collision time differences $\vartheta(\varepsilon)$, and
  6. (6)
    $\textnormal{Slow}(\varepsilon)\leq CM(\varepsilon)(J(\varepsilon))^{3}$ is to rule out collisions with relative speeds smaller than $\iota(\varepsilon)$,
@@
$+P_{t}\big(\mathcal{V}(\Phi)>V(\varepsilon)\ \mathrm{with}\ n(\Phi)\leq M(\varepsilon)\big)$
$+P_{t}\big(\vartheta(\Phi)<I(\varepsilon)\ \mathrm{with}\ n(\Phi)\leq M(\varepsilon)\big)$
$+P_{t}\big(\iota(\Phi)<J(\varepsilon)\ \mathrm{with}\ n(\Phi)\leq M(\varepsilon)\big).$
The forth and fifth terms can be directly estimated by Lemmas 4.6 and 4.7 respectively. We also observe that the third term can be written as
@@
Now, by the Markov’s inequality and by (4.2) we have
for all $t\geq 0$ and $n\in(3,5]$. Also, by Proposition 4.4 we have
@@
by using Proposition 4.2 in the last inequality, since $\tilde{M}_{f}(t)\leq D_{f}(t)$. Therefore, by summing (4.14), (4.22), (4.4) and (4.24) and the estimates in Lemmas 4.6 and 4.7, we obtain the desired estimate
@@
[…]
$\quad+CM(\varepsilon)\left(I(\varepsilon)+(J(\varepsilon))^{3}\right)$
$\quad+\frac{M_{f_{0}}}{V(\varepsilon)}+(C_{g}+\tilde{C}_{q,n}ct)\frac{M(\varepsilon)}{V(\varepsilon)},\ t\in[0,T]\ \textrm{and}\ n\in(3,5].$
@@ 4.8. The Empirical Distribution
On the set of good histories the particle dynamics can be stated in a similar way to (2.1). We consider now the empirical distribution on collision histories $\hat{P_{t}^{\varepsilon}}$ defined by the dynamics of the particle system for particles with diameter $\varepsilon$. We will write $\hat{P_{t}}$ instead of $\hat{P_{t}^{\varepsilon}}$. The main result in this subsection is that $\hat{P_{t}}$ solves a differential equation which is similar to the idealised equation (2.1) when restricted to the set of good trees $\mathcal{G}(\varepsilon)$. We follow the approach in [MST18] which derived these equations for hard-sphere interactions and adaptions made for finite and infinite range interactions in [ET17]. As initial conditions are obtained from a Poisson process the equations simplify as the terms due to conditioning do not appear.
@@
$\hat{\mathcal{Q}}^{+}_{t}[\hat{P_{t}}](\Phi):=\begin{cases}\delta(t-\tau)\hat{P_{t}}(\bar{\Phi})g_{0}(v^{\prime})\mathcal{B}(v(\tau^{-})-v^{\prime},\omega)&\text{if}\ n(\Phi)\geq 1\\ 0&\text{if}\ n(\Phi)=0.\end{cases}$
@@
$\hat{\mathcal{Q}}^{-}_{t}[\hat{P_{t}}](\Phi):=\mathds{1}_{t\geq\tau(\Phi)+T^{*}(\Phi)}\hat{P_{t}}(\Phi)\int_{\mathbb{S}^{2}}\int_{\mathds{R}^{3}}g_{0}(\bar{v})\mathcal{B}(v(\tau)-\bar{v},\omega)\mathds{1}^{\varepsilon}_{t}[\Phi](x(t)+\varepsilon\omega,\bar{v})\,\mathrm{d}\bar{v}\,\mathrm{d}\omega.$
@@
$\begin{cases}\partial_{t}{\hat{P_{t}}}(\Phi)&=c\hat{\mathcal{Q}_{t}}[\hat{P_{t}}](\Phi)\\ \hat{P_{0}}(\Phi)&=\zeta(\varepsilon)f_{0}(x_{0},v_{0})\mathds{1}_{n(\Phi)=0}\end{cases}$ (4.26)
with a correction factor for initial overlaps
$\zeta(\varepsilon):=\exp(-\tfrac{4\pi}{3}N\varepsilon^{3})$ (4.27)
due to the spatially Poisson distributed initial data.
This is a direct adaption of [MST18, Theorem 4.6.] and [Sto17] for the version of Rayleigh in (1.2). The same approach was also used in [MS18, ET17, Egg18]. Following [GSRT13, Prop 8.2.1], the collision intervals for the empiric can be bounded by $C\varepsilon(\mathcal{V}(\Phi)+\frac{1}{\iota(\Phi)})\leq C\varepsilon(V(\varepsilon)+\frac{1}{J(\varepsilon)})$, if this bound is considerably smaller than $\vartheta(\Phi)$, e.g. if
$C\varepsilon(V(\varepsilon)+\frac{1}{J(\varepsilon)})\ll I(\varepsilon),$ (4.28)
then $\Phi$ will only undergo binary collisions.
There are two substantial simplification compared to [MST18]: firstly due to the Poisson set-up we do not have terms due to conditioning on the history or due to a reduction of available particles and similarly initial overlaps can be calculated directly. In the next steps we use that the evolution defined by (4.26) is linear in $f_{0}$. We first rescale with the factor $\zeta(\varepsilon)^{-1}$ and get a simple comparison between $P_{t}(\Phi)$ and the rescaled $\hat{R}_{t}(\Phi):=\zeta(\varepsilon)^{-1}\hat{P_{t}}(\Phi)$
@@
$\hat{\mathcal{Q}}^{-}_{t}[\hat{R_{t}}](\Phi)$ $=\hat{L}_{t}(\Phi)\hat{R_{t}}(\Phi),$
$\mathcal{Q}^{-}_{t}[P_{t}](\Phi)$ $=L_{t}(\Phi)P_{t}(\Phi),$
where
$\hat{L}_{t}(\Phi):=\mathds{1}_{t\geq\tau(\Phi)+T^{*}(\Phi)}\int_{\mathbb{S}^{2}}\int_{\mathds{R}^{3}}g_{0}(\bar{v})\mathcal{B}(v(\tau)-\bar{v},\omega)\mathds{1}^{\varepsilon}_{t}[\Phi](x(t)+\varepsilon\omega,\bar{v})\,\mathrm{d}\bar{v}\,\mathrm{d}\omega$
and
$L_{t}(\Phi):=\int_{\mathds{R}^{3}}\int_{\mathbb{S}^{2}}g_{0}(\bar{v})\mathcal{B}(v^{\varepsilon}(\tau)-\bar{v},\omega)\,\mathrm{d}{\omega}\,\mathrm{d}\bar{v},$
@@
Then for a fixed $\Phi$ and $t>\tau(\Phi)$ we obtain a scalar ordinary differential inequality for $\hat{R_{t}}(\Phi)-P_{t}(\Phi)$
$\partial_{t}(\hat{R_{t}}(\Phi)-P_{t}(\Phi))$ $=-c\hat{L}_{t}(\Phi)\hat{R_{t}}(\Phi)+cL_{t}(\Phi)P_{t}(\Phi)$
$=-cL_{t}(\Phi)(\hat{R_{t}}(\Phi)-P_{t}(\Phi))+c[L_{t}(\Phi)-\hat{L}_{t}(\Phi)]\hat{R_{t}}(\Phi)$
$\geq-cL_{t}(\Phi)(\hat{R_{t}}(\Phi)-P_{t}(\Phi)).$
@@
$\hat{R_{t}}(\Phi)-P_{t}(\Phi)\geq\exp\left(-c\int_{\tau}^{t}L_{s}(\Phi)\mathrm{d}s\right)(\hat{R_{\tau}}(\Phi)-P_{\tau}(\Phi)).$ (4.31)
So we obtain (4.29) if $\hat{R_{\tau}}(\Phi)-P_{\tau}(\Phi)\geq 0$. The $n(\Phi)=0$ holds for $\tau=0$ due to the initial conditions in (2.1) and (4.26). For all other good trees this follows by induction over $n(\Phi)$ with (4.30) as the gain operator $\mathcal{Q}^{+}_{t}[P](\Phi)$ is monotone in $P$. ∎
@@
[…]
$\quad+\frac{M_{f_{0}}}{V(\varepsilon)}+(C_{g}+Cct)\frac{M(\varepsilon)}{V(\varepsilon)}$
$\quad+CM(\varepsilon)(I(\varepsilon)+(J(\varepsilon)^{3})+(1-\zeta(\varepsilon)).$ (4.32)